# From Actions to Words: Towards Abstractive-Textual Policy Summarization in RL

### v2

@@ Appendix A HIGHLIGHTS
Our user study uses “Highlights” policy summaries [17] as a baseline. For completeness, we describe the algorithm here. The HIGHLIGHTS algorithm generates an online summary of an agent’s behavior from simulations, using state importance to decide which states to include. A state is considered important if taking a wrong action there significantly decreases future rewards, as determined by the agent’s Q-values. Formally, state importance $I(s)$ is defined as:
@@ Appendix B Information Stored in Experience dataset
Table 7: Description of the data stored in the experience dataset.
@@ F.1 MiniGrid
Table 8: Hyper-parameters for the PPO algorithm applied to all six agents.
F.2 Crafter
@@ Appendix I Systematic Exploration of the Prompt Design
The creation of the final prompt was achieved through a structured and iterative exploration process. This process involved a quantitative evaluation of prompt designs based on observed outputs, guided by principles from prompt engineering literature, and tailored to domain-specific requirements. Additionally, the final design was inspired by the Chain of Thought (CoT) [47] prompting paradigm, which encourages models to generate structured, step-by-step reasoning. Below is a detailed breakdown of the methodology used:

### v3

@@ Appendix A HIGHLIGHTS
Our user study uses “Highlights” policy summaries amir2018highlights as a baseline. For completeness, we describe the algorithm here. The HIGHLIGHTS algorithm generates an online summary of an agent’s behavior from simulations, using state importance to decide which states to include. A state is considered important if taking a wrong action there significantly decreases future rewards, as determined by the agent’s Q-values. Formally, state importance $I(s)$ is defined as:
@@ Appendix B Information Stored in Experience dataset
Table 5. Description of the data stored in the experience dataset.
@@ F.1. MiniGrid
Table 6. Hyper-parameters for the PPO algorithm applied to all six agents.
F.2. Crafter
@@ Appendix I Systematic Exploration of the Prompt Design
The creation of the final prompt was achieved through a structured and iterative exploration process. This process involved a quantitative evaluation of prompt designs based on observed outputs, guided by principles from prompt engineering literature, and tailored to domain-specific requirements. Additionally, the final design was inspired by the Chain of Thought (CoT) wei2022chain prompting paradigm, which encourages models to generate structured, step-by-step reasoning. Below is a detailed breakdown of the methodology used: